OpenStudy (anonymous):

i can't figure out this problem for the life of me. it's let the region be bounded by y=x^2, x=1, and y=0. find the volume of the region about the line y=-1

OpenStudy (anonymous):

Volume, or area?

OpenStudy (anonymous):

volume

OpenStudy (anonymous):

:( I can't figure it out either

OpenStudy (anonymous):

is anything unclear

OpenStudy (anonymous):

are you rotating it?

OpenStudy (anonymous):

Yeah if it's a rotation about y = -1 then it's doable

OpenStudy (anonymous):

yes you have to rotate the region bounded by the equations around the line y=-1

OpenStudy (anonymous):

Why didn't you say so! :-P

OpenStudy (anonymous):

it made sense to me, sorry

OpenStudy (anonymous):



OpenStudy (anonymous):

that's what i thought too but the program i'm using wants it in terms of y so y=x^2 should be x=sqrt(y) so then the graph should look like this.....

OpenStudy (anonymous):